OpenStudy (anonymous):

need someone to walk me though again lol Complete the equation of the graphed linear function in point-slope form. y – (–2) = (x – ) The graph of the function goes through point 1, negative 2 and point 2, 2.

OpenStudy (anonymous):

@Ashleyisakitty

OpenStudy (anonymous):

@AnswerMyQuestions @iGreen

OpenStudy (anonymous):

@roz_mar23

OpenStudy (anonymous):

sorry..i keep dragging you into this lol

OpenStudy (anonymous):

2-(-2)/2-1

OpenStudy (igreen):

Yep^

OpenStudy (igreen):

That's the correct way to find the slope, simplify it.

OpenStudy (anonymous):

4/1

OpenStudy (igreen):

Yep, or 4.

OpenStudy (anonymous):

so slope is 4

OpenStudy (igreen):

Now plug in (1, -2) and 4 into point-slope form. \(y - y_1 = m(x - x_1)\) Where \(y_1\) is the y-value of the point, \(x_1\) is the x-value of the point, and \(m\) is the slope.

OpenStudy (anonymous):

y-(-2)=4(x-1) correct?

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

Yes!

OpenStudy (igreen):

Or: y + 2 = 4(x - 1)

OpenStudy (igreen):

\(\Huge\color{lime}\checkmark\)

OpenStudy (anonymous):

i think im getting the hang of this :)

OpenStudy (igreen):

I think so too!

OpenStudy (anonymous):

thanks for your help

OpenStudy (anonymous):

@iGreen i got a 100 on the test!